# DynVFX: Augmenting Real Videos with Dynamic Content

### v1

@@ 4. Method
Given an input video $\mathcal{V}_{\text{orig}}$ and a textual instruction $\mathcal{P}_{\text{VFX}}$, our goal is to synthesize a new video $\mathcal{V}_{\text{VFX}}$, in which new dynamic elements are seamlessly integrated to the existing scene. We address this task by estimating a residual in the latent space of the text-to-video diffusion model, which is added to the latent of the original video. The final edited video is obtained by decoding their sum with the text-to-video diffusion model’s VAE decoder
Tackling this task requires ensuring that the generated content, such as new objects or effects, adheres to the dynamics of the existing scene. The location and size of the new content must align with the camera motion and the environment, while its actions and movements must appropriately respond to other dynamic objects present in the scene. Our framework (illustrated in Fig. 3) addresses these challenges by incorporating the following key components:
@@
    VLM as a VFX assistant. We utilize a pre-trained VLM to interpret the user’s instructions, reason about the interactions with the scene’s dynamics, and provide a descriptive prompt for the T2V diffusion model by guiding it to act as ”VFX assistant” via a system prompt containing guidelines in the context of our tasks.
@@
    Localization via Anchor Extended Attention. To steer the localization of the edit and make it content-aware, we propose to utilize Extended Attention during sampling from the T2V DiT model to a set of keys and values extracted from sparse locations in the original video.
@@
    Content Harmonization via Iterative refinement. To improve the blending of the generated content with the input video and achieve better harmonization, we iteratively update the estimated residual latent by repeating the sampling process with AnchorExtAttn multiple times, progressively reducing the level of noise added at each step.
@@ 4.1. VLM as a VFX assistant
To create a fully automatic framework requiring only a simple user-provided instruction describing the desired content, we incorporate a Vision-Language Model (VLM) into our framework. Specifically, given the user instruction $\mathcal{P}_{\text{VFX}}$, along with evenly spaced keyframes from the original video $\mathcal{V}_{\text{orig}}$, we instruct the VLM (Achiam et al., 2023) to provide a detailed caption - a composition prompt $\mathcal{P}_{\text{comp}}$ describing the new composited scene. While the model gives an accurate, descriptive source scene caption, in some cases, we observed that it fails to give captions suitable for compositing VFX with the scene. To overcome this, we guide the model to reply in an in-context matter by asking it to imagine a conversation with a visual effects (VFX) artist to obtain a caption that would describe the composited scene correctly. In this conversation, VLM will “consult” with a VFX artist about how the new content should be integrated into the scene. Based on their discussion, it provides a caption that describes how the added content fits into the scene. This results in text prompts that encourage the generated output video to include a natural interaction between the new content and the original environment. The guidelines include focusing on (1) spatial and dynamic awareness of existing scene elements, (2) preservation of original scene behaviors, and (3) atmospheric coherence between new and existing content. See SM for more details about how we use the VLM to reason about the new scene integration with in-context reasoning for VFX.
We also use the VLM to obtain a list of prominent foreground objects in the original video: $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt: $\mathcal{O}_{\text{edit}}$. See more details about prompting and mask processing in SM.
@@ 4.2. Localization via Anchor Extended Attention
A pivotal aspect of our method is the accurate localization of the new content in the existing scene, ensuring alignment with camera motion, occlusions, and scene depth. While the composition prompt can describe the desired location, naive noising-denoising with this prompt introduces a trade-off: As shown in Fig. 2, using SDEdit (Meng et al., 2022) with a high noising timestep fails to retain the original scene, resulting in misaligned new content. In contrast, a low noising timestep, as illustrated in Fig. 2 (b), limits deviations from the original video.
Figure 4. Ablations. (b) Excluding both AnchorExtAttn and the Iterative refinement process results in significant misalignment with the original scene and poor harmonization (e.g., the size of the puppy relative to the scene and boundary artifacts). (c) Omitting AnchorExtAttn leads to incorrect positioning of the new content. (d) Removing iterative refinement results in poor harmonization. Our full method (e) exhibits good localization and harmonization of the edit
To tackle the localization challenge, we extend the attention module during sampling to include the input video’s corresponding attention features. Specifically, we apply DDIM inversion (Song et al., 2020) to the original video $\mathcal{V}_{\text{orig}}$ and extract the spatio-temporal keys and values $K_{\text{orig}},V_{\text{orig}}$ from the attention module of every block in the network and generation timestep $t$. These keys and values are then used to extend the attention mechanism during sampling with $\mathcal{P}_{comp}$ to control the localization of the edit.
@@
(2) $\texttt{Attn}(Q_{\text{VFX}},[K_{\text{VFX}},K_{\text{orig}}],[V_{\text{VFX}},V_{\text{orig}}])$
Interestingly, extending keys and values provides more than just global information, but each feature locally encodes corresponding patches in the video. As shown in Fig. 2 (c), extending the attention to the full set of keys and values is sufficient to achieve an approximate reconstruction of the original video. We hypothesize that this occurs because the same positional embedding is applied to the $K_{\text{orig}}$ as to $K_{\text{VFX}}$, aligning the positional embeddings in spatiotemporal locations. This observation provides strong evidence for the significance of the positional embedding. Due to this, the setup proves to be overly restrictive in adhering to $\mathcal{P}_{\text{comp}}$, as illustrated in Fig. 2(c).
@@
Selective Attention. To overcome this limitation, we propose to restrict the extended attention only to specific positions in the source scene. Specifically, we use a selection function $\mathcal{F}$ to determine which keys and values are retained.
An intuitive strategy is to apply region-based attention with Masked Extended Attention by identifying the most critical regions for preserving scene coherence and extending attention with keys and values within these masked regions $M_{\text{orig}}$, i.e., $\mathcal{F}(\mathcal{A})=M_{\text{orig}}\circ\mathcal{A}$. To get $M_{\text{orig}}$, we ask a VLM to provide a list of foreground objects in the original video $\mathcal{O}_{\text{orig}}$ (typically the most spatially prominent elements within a scene) and then we obtain corresponding masks using a text-based segmentation model. For example, Fig. 2(d), where the masked region includes the horse, illustrates the effect of this on the generated content. As shown, this setup preserves high fidelity to the original scene within the mask $M_{\text{orig}}$ and allows new content to emerge in the unmasked regions.
However, this approach proves too constrained in overlapping regions. To address this limitation, we propose our Anchor Extended Attention:
[…]
:=$\mathcal{F}:=\Biggl{\{}\texttt{Drop}_{FG}(M_{\text{orig}})\cup\texttt{Drop}_{BG}(\sim M_{\text{orig}})~{}\circ~{}\mathcal{A}$
@@
This formulation introduces dropout within the masked regions $M_{\text{orig}}$, generating a sparse set of anchor points that guide the generation while preserving flexibility. This formulation introduces a predominantly content-aware selection of anchors from foreground regions, along with a sparser set of background anchors, to achieve robust and spatially coherent integration of new content (e.g. the horse to enforce the knewly added knight to align with its motion). As demonstrated in Fig. 2(e), this balanced approach offers sufficient flexibility for creative edits while preserving key spatial cues from the original scene.
Figure 5. Sample results of our method. See SM for full vide results.
@@ 4.3. Content Harmonization
Our anchor extended attention steers the placement of the new content to align with the original scene. However, it does not guarantee precise pixel-level alignment. As can be seen in Fig. 2(e), the legs of the horse are not perfectly aligned with the original horse. To guarantee pixel-level alignment, a straightforward approach is to extract a mask of the new content $\bm{M}_{VFX}^{rgb}$ from the sampling with AnchorExtAttn (Eq. 3) output $\mathcal{\hat{V}}_{comp}$. More concretely, by applying a text-based segmentation model using the added object description provided by the VLM.
The mask can then be used to replace the pixels outside it with the corresponding pixels from the original video: $\mathcal{V}{\text{comp}}[\sim\bm{M}_{VFX}^{rgb}]=\mathcal{V}{\text{orig}}[\sim\bm{M}_{VFX}^{rgb}]$. While this preserves the unaffected regions, it often results in poor harmonization with the input video.
@@
To improve content harmonization, we propose a different approach: repeat the sampling process with AnchorExtAttn (Eq. 3) multiple times, progressively reducing the level of noise added at each step. This iterative approach gradually refines the new content’s interaction with the original scene. As shown in Fig.3, we update $\bm{x}_{res}$ representing the difference between the generated output $\mathcal{\hat{V}}_{comp}$ and the original video $\mathcal{{V}}_{orig}$ within the regions where new content appears, allowing each iteration to adjust the generated content’s high-frequency details to better match the original video. We summarize our method in Alg. 1.

### v2

@@ 4. Method
Given an input video $\mathcal{V}_{\text{orig}}$ and a textual instruction $\mathcal{P}_{\text{VFX}}$, our goal is to synthesize a new video $\mathcal{V}_{\text{VFX}}$ in which new dynamic elements are seamlessly integrated into the existing scene. Tackling this task requires ensuring that the location and size of the new content align with the camera motion and the environment, while its actions and movements must respond appropriately to other dynamic objects present in the scene. Our framework (Fig. 2) addresses these challenges by incorporating the following key components:
@@
    VLM as a VFX Assistant. We utilize a pre-trained VLM to interpret the user’s instructions, reason about the interactions with the scene’s dynamics, and identify both the prominent existing elements in the scene and the new content to be added. The VLM provides a descriptive prompt of the desired scene augmentation, a list of prominent existing elements, and the new content to be added. We achieve this by guiding the VLM to act as a “VFX assistant” via a system prompt containing guidelines in the context of our tasks.
@@
    Localization via Anchor Extended Attention. Our approach achieves precise placement of new content by guiding the T2V generation model during sampling to be content-aware. We steer the localization of the edit by applying Extended Attention with a sparse set of spatiotemporal locations extracted from the original scene, allowing the generation to focus on essential elements. Our approach is based on our observation that in the pre-trained T2V DiT model, keys and values locally encode corresponding video patches in both space and time.
@@
    Content Harmonization via Iterative Refinement. To guarantee precise pixel-level alignment of the generated content with the input video and achieve better harmonization, we iteratively refine the estimated edit by repeating the sampling process with AnchorExtAttn multiple times, while preserving all unedited regions.
Our method does not require user-provided masks to pre-determine the location of the new content, but rather infers it automatically. We leverage a text-based segmentation model (Zhang et al., 2024a) for two purposes: (i) Localization, by segmenting prominent elements from the original scene to apply AnchorExtAttn, which allows the T2V model to infer the natural location and dynamics of the new content based on the original scene context. (ii) Harmonization, by segmenting the newly added content to ensure pixel-level alignment with the original scene.
@@ 4.1. VLM as a VFX Assistant
To create a fully automatic framework that requires only a simple user instruction, we incorporate a VLM into our framework. Specifically, given a user instruction $\mathcal{P}_{\text{VFX}}$, along with keyframes from $\mathcal{V}_{\text{orig}}$, we instruct the VLM (Achiam et al., 2023) to provide a composition prompt $\mathcal{P}_{\text{comp}}$ describing the newly augmented scene. We use $\mathcal{P}_{\text{VFX}}$ as the edit prompt for the T2V generation process. Hence, the edit prompt must include the following details: (i) original scene elements and actions, (ii) new content elements and actions, and (iii) the interaction between them. When simply asking the VLM to yield $\mathcal{P}_{\text{VFX}}$, it often fails to produce captions that are both true to the original video and descriptive of the new content well. Therefore, using these for T2V generation can result in edits with low fidelity to the instruction (see examples in Sec. E.2 and Fig. 15 in the SM).
To resolve this issue, we guide the VLM in an in-context manner by instructing it via a system prompt to imagine a conversation with a VFX artist to obtain a caption that would describe the composited scene correctly. We observe that this helps the VLM to achieve better scene understanding. Specifically, the system prompt includes guidelines to focus on: (1) spatial and dynamic awareness of existing scene elements, (2) preservation of original scene behaviors, and (3) atmospheric coherence between new and existing content.
Additionally, we utilize the VLM to yield a list of prominent foreground objects in the original video $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt $\mathcal{O}_{\text{edit}}$. See additional details in Sec. A.3, Sec. A.4, and Sec. E in the SM. In addition, Figs. 16-19 in the SM include an overview of the protocol, system prompts, and output examples.
@@ 4.2. Localization via Anchor Extended Attention
A pivotal aspect of our method is the accurate placement of the new content. While $\mathcal{P}_{\text{comp}}$ can describe the desired location, naive noising-denoising with this prompt introduces a trade-off: As shown in Figs. 3 (a)-3(b), using SDEdit (Meng et al., 2022) with a high noising timestep fails to retain the original scene, resulting in misaligned new content, whereas a low noising timestep limits deviations from the original video.
To tackle the localization challenge, we extend the attention module during sampling to include the input video’s corresponding attention features. Specifically, we apply DDIM inversion (Song et al., 2020) to the original video $\mathcal{V}_{\text{orig}}$ and extract the spatiotemporal keys and values $K_{\text{orig}},V_{\text{orig}}$ from the attention module of every block in the network and generation timestep $t$. These keys and values are then used to extend the attention mechanism during sampling with $\mathcal{P}_{comp}$ to control the localization of the edit.
@@
(2) $\texttt{Attn}(Q_{\text{VFX}},[K_{\text{VFX}},K_{\text{orig}}],[V_{\text{VFX}},V_{\text{orig}}])\,.$
We observe that extending attention to the full set of keys and values approximately reconstructs the original video (Fig. 3(c)). This demonstrates that the keys and values of attention layers in the DiT-based model determine more than just appearance information. In contrast to UNet-based architectures, keys and values locally encode corresponding video patches. We hypothesize that this occurs because the same positional embedding is applied to $K_{\text{orig}}$ as to $K_{\text{VFX}}$. Remarkably, applying the same positional embedding to both source and target keys and values enables alignment between the generated and original content at corresponding spatiotemporal locations.
Figure 4. Ablations. (b) Excluding both AnchorExtAttn and the Iterative refinement process results in significant misalignment with the original scene and poor harmonization (e.g., the size of the puppy relative to the scene and boundary artifacts). (c) Omitting AnchorExtAttn leads to incorrect positioning of the new content. (d) Removing iterative refinement results in poor harmonization. Our full method (e) exhibits good localization and harmonization of the edit.
@@
Selective Attention. Based on our observation, we propose restricting the extended attention only to specific positions in the source scene. Specifically, we use a selection function $\mathcal{F}$ to determine which keys and values are retained.
Masked Extended Attention. A straightforward choice is region‐based masking: identifying the most critical regions for preserving scene coherence and extending attention with keys and values within these masked regions $M_{\text{orig}}$, i.e., $\mathcal{F}(x)=M_{\text{orig}}\circ x$.
To get $M_{\text{orig}}$, we ask a VLM for $\mathcal{O}_{\text{orig}}$, a list of foreground elements in $\mathcal{V}_{\text{orig}}$ (typically the most spatially prominent elements within a scene), and obtain the corresponding masks using a text-based segmentation model (see Sec. A.3 and Sec. A.4 in the SM for more details and examples of $M_{\text{orig}}$).
As shown in Fig. 3(d), this strategy successfully preserves high fidelity to the original scene within $M_{\text{orig}}$ while allowing new content to emerge in unmasked regions. However, Fig. 3(d) reveals two issues: (i) due to the hard constraint in overlapping regions, it is not suitable for handling natural occlusions, and (ii) the lack of guidance in unmasked regions can lead to inaccurate scale or positioning relative to the input scene.
Anchor Extended Attention. To achieve robust and spatially coherent integration of new content, we propose to selectively extend keys/values in a sparse set of foreground anchors and a sparser set of background anchors. Introducing Anchor Extended Attention with its formulation:
(3) $\texttt{AnchorExtAtt}$= $\texttt{Attn}(Q_{\text{VFX}},[K_{\text{VFX}},K^{E}],[V_{\text{VFX}},V^{E}])\,$,
$\text{s.t. }K^{E}=\mathcal{F}(K_{\text{orig}})~~and~~V^{E}=\mathcal{F}(V_{\text{orig}})\,$,
$\mathcal{F}(x)=\texttt{Drop}_{\text{FG}}(M_{\text{orig}})\cup\texttt{Drop}_{\text{BG}}(\sim M_{\text{orig}})~\circ~x\,.$
@@
For each layer, a selection of anchor points is randomly sampled via dropout, in both masked and unmasked regions. To steer the localization of the edit, we extend the attention with anchors that correspond to spatiotemporal locations in $\mathcal{V}_{\text{orig}}$, allowing the generation to focus on essential elements (anchors) to achieve robust and spatially coherent integration of new content. (e.g., the knight’s motion is aligned with the horse in Fig. 3(e)). This balanced approach offers flexibility for creative edits while preserving key spatial cues from the original scene.
@@ 4.3. Content Harmonization
Our anchor extended attention steers the placement of the new content to align with the original scene. However, it does not guarantee precise pixel-level alignment. As can be seen in Fig. 3(e), the legs of the horse are not perfectly aligned with the original video. To guarantee pixel-level alignment, a straightforward approach is to extract a mask of the new content $\bm{M}_{\text{VFX}}^{\text{rgb}}$ from the sampling with AnchorExtAttn (Eq. 3) output $\mathcal{\hat{V}}_{\text{comp}}$. More concretely, we obtain this mask by applying a text-based segmentation model using the added object description provided by the VLM. The mask can then be used to replace the pixels outside it with the corresponding pixels from the original video: $\mathcal{V}{\text{comp}}[\sim\bm{M}_{\text{VFX}}^{\text{rgb}}]=\mathcal{V}{\text{orig}}[\sim\bm{M}_{\text{VFX}}^{\text{rgb}}]$. While this preserves the unaffected regions, it often results in poor harmonization with the input video.
@@
To improve content harmonization, we propose repeating the sampling process with AnchorExtAttn (Eq. 3) multiple times, progressively reducing the level of noise added at each step. This iterative approach gradually refines the new content’s interaction with the original scene.
As shown in Fig. 2, we update $\bm{x_{\text{comp}}}$ by adding a residual latent $\bm{x}_{\text{res}}=\bm{{M}_{\text{VFX}}\cdot(\hat{x}_{\text{comp}}-x_{\text{orig}}})$ to $\bm{x_{\text{orig}}}$, where $\bm{{M}_{{\text{VFX}}}}$ is the latent representation of $\bm{{M}_{\text{VFX}}^{\text{rgb}}}$ (see Sec. A.4 in SM for details on computing the latent $\bm{{M}_{\text{VFX}}}$). Notably, the addition of $\bm{x}_{\text{res}}$ is equivalent to directly replacing values of $\bm{x_{\text{orig}}}$ with those of $\bm{\hat{x}_{\text{comp}}}$ within $\bm{{M}_{\text{VFX}}}$. The final edited video is obtained by decoding their sum with the T2V model’s VAE decoder. This allows each iteration to adjust the generated content’s high-frequency details to better match the original video. We provide implementation details and summarize our method in Sec. A and Alg. 1 in the SM.
Figure 5. Sample Results of DynVFX. Our method supports a wide range of scene augmentations across diverse scenarios while maintaining realistic interaction, occlusion, lighting, and camera motion, for example: a golden retriever consistent with camera movement, transparent wings revealing the woman’s silhouette at sunset, and a tsunami flooding the city yet realistically respecting the car dashboard. See SM for full videos.
Figure 6. Qualitative Comparison of Text-Based Methods. Sample results comparing our method to SDEdit (Meng et al., 2022), DDIM inversion (Song et al., 2020), Lora fine-tuning (Hu et al., 2022), Gen-3 (Runway, 2024) and FlowEdit (Kulikov et al., 2024). As can be seen, our method better augments the original scene with new dynamic content that interacts naturally with existing elements in the scene. See SM for full video comparison.
5. Results
We evaluated our method over a set of 57 video-text edit pairs containing 36 unique publicly available videos from the web and the DAVIS dataset (Pont-Tuset et al., 2017). These videos feature a wide range of complex scenes in terms of camera and object motion, lighting conditions, and physical environments. Our videos and implementation details are available in the SM.
Figs. 1 and  5 show sample results of our method. As seen, our method facilitates natural integration of a broad range of visual effects, ranging from environmental effects (a tsunami in Fig. 4 and an explosion in Fig. 5) to new object insertion (knight riding a horse in Fig. 5 and dancing bear in Fig. 6). In all examples, the new content is naturally localized in the scene, even in challenging scenarios involving multiple objects (dinosaurs or workers in Fig. 5) and partial occlusions (a puppy in Fig. 1 and a giraffe in Fig. 5).
Figure 7. Comparison to Two-Stage Editing Methods. Left pane: our results. Right pane: AnyV2V and I2VEdit. See SM for video comparisons.
5.1. Qualitative Evaluation
We compare our method to the following baselines: (i) SDEdit (Meng et al., 2022) using the same T2V model as ours, (ii) DDIM inversion (Song et al., 2020) and sampling with the target prompt, (iii) LORA fine-tuning (Hu et al., 2022) of the T2V model and sampling with the target prompt, (iv) Gen-3 (Runway, 2024) video-to-video model, designed for video stylization, (v) FlowEdit (Kulikov et al., 2024) and two-stage editing approaches: (vi) AnyV2V (Ku et al., 2024) and (vii) I2VEdit (Ouyang et al., 2024).
Figs. 6 and 7 show a qualitative comparison with the baselines. As can be seen, all baselines exhibit different limitations in maintaining scene fidelity while introducing new content. SDEdit (Meng et al., 2022) and FlowEdit (Kulikov et al., 2024) both manage to fulfill the edit prompt, yet the scene may significantly deviate from the original scene in terms of appearance, motion, positioning, or scale (e.g., deer in the creek). DDIM inversion is not suitable for editing. LORA fine-tuning suffers from the trade-off between preserving aspects of the original scene and adding new content to the scene: either overfitting original scene appearance (e.g., adding a dragon-dog hybrid), or underfitting the original layout (e.g., incorrect scale of deer). Gen-3 is conditioned on structure and content representation extracted from the input video, hence it tends to significantly alter scene appearance and does not allow the insertion of new objects that change the scene layout. Both AnyV2V and I2VEdit take the first edited frame as additional input, but still struggle to maintain the coherent motion of the generated objects. In particular, AnyV2V often fails to accurately propagate the edits through the video, leading to distorted geometries. In each case, these limitations affect the overall scene coherence and realism of the added elements. Our method successfully adds new content to the scene, achieving high fidelity to the user instructions while allowing for natural interactions between original and added elements (e.g., natural interaction between woman and bear). We provide details about the baselines comparison in Sec. B in the SM. Additionally, we provide qualitative comparisons with VFX reference-based methods (Pika, 2024; KlingAI, 2024) in Sec. D and Fig. 12 in the SM.
Table 1. Quantitative Comparison. We compare our method with other video editing methods and ablations using metrics for edit fidelity (CLIP Directional) and original content preservation (masked SSIM), a VLM-based evaluation of text alignment, visual quality, harmonization, and motion realism, and a user study on content integration and edit harmonization. User study values indicate the percentage of participants who preferred our method over each baseline.
Method Metrics VLM-based evaluation User Study
Gen-3 0.142 0.283 0.410 0.586 0.363 0.374 98.687 94.691
LORA finetuning 0.292 0.349 0.781 0.755 0.726 0.714 92.221 71.968
DDIM inv. sampling 0.193 0.440 0.465 0.609 0.448 0.448 98.951 93.145
SDEdit (0.9) 0.290 0.321 0.778 0.750 0.717 0.730 99.154 75.645
SDEdit (0.6) 0.105 0.558 0.381 0.703 0.406 0.390 92.655 92.470
FlowEdit 0.288 0.414 0.771 0.765 0.712 0.736 97.346 65.494
AnyV2V 0.300 0.468 0.678 0.669 0.615 0.619 91.142 77.747
I2VEdit 0.288 0.524 0.769 0.773 0.724 0.719 91.030 82.563
w/o AnchorExtAttn 0.325 0.691 0.765 0.713 0.670 0.687 81.417 81.109
w/o Iterative Refinement 0.291 0.736 0.749 0.741 0.697 0.696 80.122 80.492
w/o VLM Protocol 0.254 0.741 0.705 0.755 0.669 0.667 71.360 73.708
Ours 0.307 0.749 0.843 0.784 0.778 0.775 - -
5.2. Quantitative Evaluation
We numerically evaluate our results using the following metrics:
(i) Edit fidelity. Following previous works (e.g., (Hsu et al., 2025; Tewel et al., 2024a)), we measure per-frame Directional CLIP Similarity (Radford et al., 2021; Gal et al., 2022) to assess the alignment between the change in the source and the target prompt, and the change between the source and edited frames.
(ii) Original content preservation. We evaluate how well the edited video preserves the original content outside the modified region. To this end, we segment the new content in the edited video using (Zhang et al., 2024a), and compute the masked Structural Similarity Index (SSIM) over the regions complementary to the edited ones.
(iii) VLM quality evaluation. Inspired by (Hsu et al., 2025; Huang et al., 2024), we employ a VLM to expand the per-frame metrics above as follows. We input several frames from the edited videos into the VLM and instruct it to evaluate four key aspects: how well the edit follows the text prompt (Text Alignment), the overall visual quality of the edited frames (Visual Quality), how well the new content is harmonized with the source frames (Edit Harmonization), and the realism of the added object’s dynamics relative to the scene (Dynamics Score). For each aspect, the VLM outputs a score between 0 and 1, with higher scores indicating better performance. In the SM, we include our evaluation protocol and the justification of its use as an evaluation metric, showing its alignment with human preferences and consistent outcomes with an alternative VLM backbone. We include this in Sec. F and Table 2 in the SM.
Table 1 and Fig. 9 in the SM present the results of the described metrics on a set of 57 video-text edit pairs comprising 34 unique videos. As shown, our method outperforms the baselines in both SSIM and Directional CLIP metrics, demonstrating superior edit fidelity while maintaining higher structural similarity in the unedited regions. The VLM-based evaluation aligns with this assessment and further shows that our method produces videos that achieve better content integration and greater motion realism.
(iv) User study. We conducted a user study to evaluate the ability to integrate new content while preserving the original video. Participants were shown the input video, a text description of the new content, our result, and a baseline output. They were asked two questions: “Which video better preserves the original footage while adding new content?” and “Which video better integrates the new content in a realistic and seamless way?”. In total, we collected 17,160 user judgments from 330 users. As seen in Table 1, our method is consistently preferred over all baselines.
5.3. Ablations
We ablate key design choices of our method: Anchor Extended Attention, iterative updates of the edit, and our VLM as a VFX Assistant protocol, by excluding each component from our framework. The ablation of our VLM protocol can be found in Sec. E.2 and Fig. 15 in the SM.
As seen in Fig. 4(c), omitting AnchorExtAttn leads to new content being misaligned relative to the original scene, with the added content poorly integrated into the original scene. Applying only the first iteration of our method (w/o iterative refinement, Fig. 4(d)) results in a better alignment with the input video, but the composition is still unstable, as evident, for example, in the puppy’s body hovering over the box in the first scene. Our full method achieves better composition with proper spatial relationships, demonstrating the importance of both components for realistic scene editing (Fig. 4(f)). We numerically evaluate each ablation with the same set of metrics described in Sec 5.2 and report them in Table 1. An additional ablation example can be found in Fig. 14 in the SM.
Figure 8. Limitations. In some cases, the T2V model struggles to precisely follow the edit prompt.
6. Discussion and Conclusions
We introduced the task of augmenting real videos with new dynamic content based on a user-provided instruction. We presented a zero-shot method utilizing the T2V model in a feature manipulation framework, enabling correct localization and natural blending of new content with existing video elements.
As our method is built upon the pre-trained T2V model, the quality of the generated edits is inherently tied to the performance and capabilities of the underlying model. As shown in Fig. 8, the T2V model sometimes struggles to generate videos that precisely follow the edit prompt; additional failure cases and limitations are detailed in Sec. C and Fig. 11 in the SM. While the quality of the edits is bound by the capabilities of the T2V model, our method is model-agnostic and can be re-implemented with future transformer-based T2V models. We anticipate its performance will improve as more powerful models become available. Even under these constraints, our method significantly outperforms baselines, expanding the capabilities of pre-trained text-to-video diffusion models.
Appendix
Appendix A Implementation Details
A.1. Models
Text-to-Video Model. We use a publicly available CogVideoX-5B (Hong et al., 2022; Yang et al., 2024) text-to-video model, which can generate videos with up to $480\times 720$ pixels resolution, 6 seconds in length, 49 frames at 8 fps. This model is a transformer-based model that processes both text and video modalities together.
Segmentation Model. To segment the prominent objects in the video and the newly generated content, we utilize EVF-SAM2 (Zhang et al., 2024a) - a text-based video segmentation model based on SAM2 (Ravi et al., 2024).
Visual Language Model. Our vision-language model of choice is GPT-4o (Achiam et al., 2023), which we use through the official OpenAI API.
Figure 9. Metrics Visualization. We measure CLIP Directional score (higher is better) and masked SSIM (higher is better). Our method demonstrates a better balance between these two metrics. See Sec. 5.2 for full discussion.